OpenStudy (anonymous):

Need help understanding this problem and how to get this answer. Please explain. Quadratic Equations w^4-4w^2-2=0 Answer: sqrt{2+sqrt{6}}, sqrt{2-sqrt{6} -sqrt{2+sqrt{6}}, -sqrt{2-sqrt{6}}

OpenStudy (anonymous):

\[\sqrt{2+\sqrt{6}} , \sqrt{2-\sqrt{6}}\]

OpenStudy (anonymous):

\[-\sqrt{2+\sqrt{6}} , -2\sqrt{-\sqrt{6}}\]

OpenStudy (anonymous):

2 should be inside the radical my bad

OpenStudy (mertsj):

I think we might want to say: \[w^2=x\]

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

Then our equation becomes: \[x^2-4x-2=0\]

OpenStudy (anonymous):

my book says that w^2 = u

OpenStudy (mertsj):

And we can use the quadratic formula on that.

OpenStudy (mertsj):

Ok. Then use u

OpenStudy (mertsj):

Then when we find the two values of u, we can replace u with w^2 and solve those two equations.

OpenStudy (anonymous):

im still lost, i understand up to that but than they get these radicals out of no where. What do i multiply add divide or what not

OpenStudy (anonymous):

do I use -b +/- √b^2-4ac/2a

OpenStudy (anonymous):

to get to that answer, unfortunately my book doesnt explain very well

OpenStudy (mertsj):

\[u=\frac{4\pm \sqrt{16-4(-2)}}{2}=\frac{4\pm \sqrt{24}}{2}=\frac{4\pm2\sqrt{6}}{2}=2\pm \sqrt{6}\]

OpenStudy (mertsj):

But u = w^2

OpenStudy (mertsj):

\[w^2=2+\sqrt{6}\]

OpenStudy (anonymous):

ah ok. I see where i went wrong. My book just skipped that whole step loll was confused that they were still referring to previous chapters of work

OpenStudy (mertsj):

\[w=\pm \sqrt{2+\sqrt{6}}\]

OpenStudy (anonymous):

yea I see now thanks a lot. Was just lost

OpenStudy (mertsj):

yw